# On the Parameter Identifiability of Partially Observed Linear Causal Models

### v1

@@ A.2 Proof of Theorem 1
Consider a graph $\mathcal{G}$ that satisfies (i) and (ii) in Theorem 1. For an atomic cover $\mathbf{V}$ in $\mathcal{G}$ with one latent variable, $\mathbf{V}=\{\mathsf{L}\}\cup\{\mathsf{X_{i}}\}_{i=1}^{m}$ ($m$ could be zero), if it has an observed pure child $\mathsf{C}$ and the coefficients of the edges from $\mathbf{V}$ to $\mathsf{C}$, i.e., $f_{\mathsf{L}\rightarrow\mathsf{C}}$, $f_{\mathsf{X_{1}}\rightarrow\mathsf{C}}$, …, $f_{\mathsf{X_{m}}\rightarrow\mathsf{C}}$, are known, then for any variable $\mathsf{A}$, such that $\mathsf{A}\neq\mathsf{C}$ and $\mathsf{A}$ is not a descendant of $\mathsf{C}$, $\sigma_{\mathsf{L},\mathsf{A}}$ can be calculated as $(\sigma_{\mathsf{A},\mathsf{C}}-\Sigma_{i=1}^{m}\sigma_{\mathsf{X_{i}},\mathsf{A}}f_{\mathsf{X_{i}}\rightarrow\mathsf{C}})/f_{\mathsf{L}\rightarrow\mathsf{C}}$.
@@
By the definition of atomic covers, all variables in $\mathbf{V}=\{\mathsf{L}\}\cup\{\mathsf{X_{i}}\}_{i=1}^{m}$ are not adjacent. By trek rule and the fact that $\mathsf{C}$ is a pure child of $\mathbf{V}$, all treks from $\mathsf{C}$ to $\mathsf{A}$ go through $\mathbf{V}$, and thus by the trek rule we have $\sigma_{\mathsf{L},\mathsf{A}}f_{\mathsf{L}\rightarrow\mathsf{C}}+\Sigma_{i=1}^{m}\sigma_{\mathsf{X_{i}},\mathsf{A}}f_{\mathsf{X_{i}}\rightarrow\mathsf{C}}=\sigma_{\mathsf{A},\mathsf{C}}$. ∎
Remark: This lemma implies that we can find all edge coefficients of the graph in a bottom-up fashion. Roughly speaking, for a latent variable $\mathsf{L}$ that belongs to an atomic cover $\mathbf{V}$, once we identify $f_{\mathsf{L}\rightarrow\mathsf{C}}$, $f_{\mathsf{X_{1}}\rightarrow\mathsf{C}}$, …, $f_{\mathsf{X_{m}}\rightarrow\mathsf{C}}$ where $\mathsf{C}$ is a pure child of $\mathbf{V}$, we can take $\mathsf{L}$ as if it is observed. More specific explanations can be found in the following proof.
See 1
@@
Consider a graph $\mathcal{G}$ that satisfies (i) and (ii) in Theorem 1. We first show that if all the pure children of an atomic cover $\mathbf{V}$ are observed, then all the edge coefficients from the atomic cover to its children are identifiable (upto group sign). To this end, we categorize the scenarios into four cases and prove them separately with illustrative examples.
@@
Example. Take the atomic cover $\{\mathsf{L_{4}}\}$ in Figure 5 as an example. By Condition 1, it must have at least two pure children. To identify $f_{4,13}$ and $f_{4,14}$, we need to borrow an observed variable that has at least one trek from $\mathsf{L_{4}}$ to it, e.g, $\mathsf{X_{8}}$. As $\mathsf{L}$ must have two other neighbors, there must exist such a variable. Then we have $\sigma_{13,14}=f_{4,13}f_{4,14}$, $\sigma_{13,8}=f_{4,13}\sigma_{4,8}$, and $\sigma_{14,8}=f_{4,14}\sigma_{4,8}$. By these three equations, we can solve $f_{4,13}$ and $f_{4,14}$.
@@
(c) $\mathbf{V}=\{\mathsf{L}\}\cup\{\mathsf{X_{i}}\}_{i=1}^{m}$ ($m\geq 1$) contains a single latent variable and $m$ observed variables, where $\mathbf{V}$ has at least three pure children. We assume that there exist $\mathsf{C_{1}}$, $\mathsf{C_{2}}$, and $\mathsf{C_{3}}$, such that $\mathsf{C_{1}}$, $\mathsf{C_{2}}$, and $\mathsf{C_{3}}$ are pure children of $\mathbf{V}$. By (ii) in Theorem 1, there exist $\mathbf{\hat{X}}$ such that $\mathbf{\hat{X}}$ d-separates $\mathsf{L}$ and $\{\mathsf{X_{i}}\}_{i=1}^{m}$. Let $\sigma_{\epsilon_{\mathsf{L}}}=t$ (the variance of the noise term of $\mathsf{L}$). In this case, we have
$\sigma_{\mathsf{C_{1}},\mathsf{C_{2}}|\mathbf{\hat{X}}\cup\{\mathsf{X_{i}}\}_{i=1}^{m}}=tf_{\mathsf{L}\rightarrow\mathsf{C_{1}}}f_{\mathsf{L}\rightarrow\mathsf{C_{2}}},$ (5)
$\quad\sigma_{\mathsf{C_{1}},\mathsf{C_{3}}|\mathbf{\hat{X}}\cup\{\mathsf{X_{i}}\}_{i=1}^{m}}=tf_{\mathsf{L}\rightarrow\mathsf{C_{1}}}f_{\mathsf{L}\rightarrow\mathsf{C_{3}}},$ (6)
$\sigma_{\mathsf{C_{2}},\mathsf{C_{3}}|\mathbf{\hat{X}}\cup\{\mathsf{X_{i}}\}_{i=1}^{m}}=tf_{\mathsf{L}\rightarrow\mathsf{C_{2}}}f_{\mathsf{L}\rightarrow\mathsf{C_{3}}}.$ (7)
@@
By these three equations, we can solve $f_{\mathsf{L}\rightarrow\mathsf{C_{1}}}$,$f_{\mathsf{L}\rightarrow\mathsf{C_{2}}}$, and $f_{\mathsf{L}\rightarrow\mathsf{C_{3}}}$, with the only remaining free parameter $t$. In other words, we have $f_{\mathsf{L}\rightarrow\mathsf{C_{1}}}(t)$,$f_{\mathsf{L}\rightarrow\mathsf{C_{2}}}(t)$, and $f_{\mathsf{L}\rightarrow\mathsf{C_{3}}}(t)$. Now we consider $\mathbf{\hat{X}}=\{\mathsf{\hat{X}_{j}}\}_{j=1}^{p}$ to solve $t$. We have additional $pm+p$ free parameters that are $f_{\mathsf{\hat{X}_{j}}\rightarrow\mathsf{X_{i}}},i\leq m,j\leq p$ and $f_{\mathsf{\hat{X}_{j}}\rightarrow\mathsf{L}},j\leq p$, and they can be solved by $\sigma_{\mathsf{\hat{X}_{j}},\mathsf{X_{i}}},~{}i\leq m,j\leq p$ and $\sigma_{\mathsf{\hat{X}_{j}},\mathsf{L}}(t),j\leq p$ (where $\sigma_{\mathsf{\hat{X}_{j}},\mathsf{L}}(t)$ is calculated by Lemma A.1). To this point, all the edge coefficients among $\mathbf{V},\mathbf{\hat{X}}$, and $\mathsf{C_{1}}$,$\mathsf{C_{2}}$,$\mathsf{C_{3}}$ are solved, upto the only free parameter $t$. Finally, we have that $\mathsf{L}$ has unit variance and thus we have $t+\sum_{j=1}^{p}f_{\mathsf{\hat{X}_{j}}\rightarrow\mathsf{L}}^{2}(t)=1$. By this additional equation, $t$ can also be solved and thus all desired edge coefficients are identified. If $\mathbf{V}$ has more than three pure children, we just choose all the combinations of any three and thus we can identify all the edge coefficients from $\mathbf{V}$ to all its pure children.
Example. Take the atomic cover $\{\mathsf{L_{1}},\mathsf{X_{5}}\}$ in Figure 6 as an example. Let $\sigma_{\epsilon_{\mathsf{L_{1}}}}=t$. We have $\sigma_{\mathsf{X_{8}},\mathsf{X_{9}}|\{\mathsf{X_{3}},\mathsf{X_{5}}\}}=tf_{\mathsf{L_{1}}\rightarrow\mathsf{X_{8}}}f_{\mathsf{L_{1}}\rightarrow\mathsf{X_{9}}}$, $\sigma_{\mathsf{X_{8}},\mathsf{X_{10}}|\{\mathsf{X_{3}},\mathsf{X_{5}}\}}=tf_{\mathsf{L_{1}}\rightarrow\mathsf{X_{8}}}f_{\mathsf{L_{1}}\rightarrow\mathsf{X_{10}}}$, and $\sigma_{\mathsf{X_{9}},\mathsf{X_{10}}|\{\mathsf{X_{3}},\mathsf{X_{5}}\}}=tf_{\mathsf{L_{1}}\rightarrow\mathsf{X_{9}}}f_{\mathsf{L_{1}}\rightarrow\mathsf{X_{10}}}$. By these three equations, we can solve $f_{\mathsf{L_{1}}\rightarrow\mathsf{X_{8}}}(t)$,$f_{\mathsf{L_{1}}\rightarrow\mathsf{X_{9}}}(t)$, and $f_{\mathsf{L_{1}}\rightarrow\mathsf{X_{10}}}(t)$, upto $t$. Then we introduce $\mathsf{X_{3}}$ and solve $f_{\mathsf{\mathsf{X_{3}}}\rightarrow\mathsf{X_{5}}}$ and $f_{\mathsf{{X}_{3}}\rightarrow\mathsf{L_{1}}}$ by $\sigma_{\mathsf{X_{3}},\mathsf{X_{5}}}$ and $\sigma_{\mathsf{X_{3}},\mathsf{L_{1}}}(t)$. Finally, we have $t+f_{\mathsf{X_{3}}\rightarrow\mathsf{L_{1}}}^{2}(t)=1$ and thus all desired edge coefficients are identified.
@@
(d) $\mathbf{V}=\{\mathsf{L}\}\cup\{\mathsf{X_{i}}\}_{i=1}^{m}$ ($m\geq 1$) contains a single latent variable and $m$ observed variables, where $\mathbf{V}$ has two pure children. By Condition 1, there must exist $\mathsf{C_{1}}$, $\mathsf{C_{2}}$ as the pure children of $\mathbf{V}$. If there exists one additional pure child, then it is the same as (c). By (ii) in Theorem 1, there exist $\mathbf{\hat{X}}=\{\mathsf{\hat{X}}_{j}\}_{j=1}^{p}$ such that $\mathbf{\hat{X}}$ d-separates $\mathsf{L}$ and $\{\mathsf{X_{i}}\}_{i=1}^{m}$. Plus, if there only exist these two pure children, there must exist $\mathsf{X_{N}}$ such that $\mathsf{X_{N}}$ is a parent of $\mathbf{V}$, and thus $\mathsf{X_{N}}\in\mathbf{\hat{X}}$. Without loss of generality, we take $\mathsf{X_{N}}=\mathsf{\hat{X}}_{1}$.
We assume that $\sigma_{\epsilon_{\mathsf{L}}}=t$, and thus we have
$\sigma_{\mathsf{C_{1}},\mathsf{C_{2}}|\mathbf{\hat{X}}\cup\{\mathsf{X_{i}}\}_{i=1}^{m}}$ $=tf_{\mathsf{L}\rightarrow\mathsf{C_{1}}}f_{\mathsf{L}\rightarrow\mathsf{C_{2}}},$ (8)
$\sigma_{\mathsf{C_{1}},\mathsf{X_{i}}|\mathbf{\hat{X}}}$ $=\sigma_{\mathsf{X_{i}}|\mathbf{\hat{X}}}f_{\mathsf{X_{i}}\rightarrow\mathsf{C_{1}}},i\leq m,$ (9)
$\sigma_{\mathsf{C_{2}},\mathsf{X_{i}}|\mathbf{\hat{X}}}$ $=\sigma_{\mathsf{X_{i}}|\mathbf{\hat{X}}}f_{\mathsf{X_{i}}\rightarrow\mathsf{C_{2}}},i\leq m.$ (10)
We further have $\sigma_{\mathsf{C_{1}},\mathsf{X_{N}}|(\mathbf{\hat{X}}\backslash\mathsf{X_{N}})}=f_{\mathsf{X_{N}}\rightarrow\mathsf{L}}f_{\mathsf{L}\rightarrow\mathsf{C_{1}}}+\sum_{i=1}^{m}f_{\mathsf{X_{N}}\rightarrow\mathsf{X_{i}}}f_{\mathsf{X_{i}}\rightarrow\mathsf{C_{1}}}$$\sigma_{\mathsf{C_{2}},\mathsf{X_{N}}|(\mathbf{\hat{X}}\backslash\mathsf{X_{N}})}=f_{\mathsf{X_{N}}\rightarrow\mathsf{L}}f_{\mathsf{L}\rightarrow\mathsf{C_{2}}}+\sum_{i=1}^{m}f_{\mathsf{X_{N}}\rightarrow\mathsf{X_{i}}}f_{\mathsf{X_{i}}\rightarrow\mathsf{C_{2}}}$,  and $\sigma_{\mathsf{X_{N}},\mathsf{X_{i}}|(\mathbf{\hat{X}}\backslash\mathsf{X_{N}})}=f_{\mathsf{X_{N}}\rightarrow\mathsf{X_{i}}},i\leq m$. We now have $33m+$ 3 edge coefficients and $33m+$ 3 equations, and by these equations, we can solve these edge coefficients upto $t$ (i.e., all these edge coefficients are functions of $t$).
Finally we solve $t$ by considering variables in $(\mathbf{\hat{X}}\backslash\mathsf{X_{N}})=\{\mathsf{\hat{X}}_{j}\}_{j=2}^{p}$. Specifically, we have $f_{\mathsf{\hat{X}}_{j}\rightarrow\mathsf{X_{i}}}=\sigma_{\mathsf{\hat{X}}_{j},\mathsf{X_{i}}|(\mathbf{\hat{X}}\backslash\mathsf{\hat{X}}_{j})}$, where $i\leq m,2\leq j\leq p$. We also have $f_{\mathsf{\hat{X}}_{j}\rightarrow\mathsf{L}}=\sigma_{\mathsf{\hat{X}}_{j},\mathsf{L}|(\mathbf{\hat{X}}\backslash\mathsf{\hat{X}}_{j})}(t)$, where $2\leq j\leq p$. By the unit variance of $\mathsf{L}$, we further have $t+\sum_{j=1}^{p}f_{\mathsf{\hat{X}_{j}}\rightarrow\mathsf{L}}^{2}(t)=1$, by which $t$ can be solved. Therefore, all edge coefficients among $\mathbf{V},\mathbf{\hat{X}},\mathsf{C_{1}},\mathsf{C_{2}}$ can be identified.
Example. Take the atomic cover $\{\mathsf{L_{2}},\mathsf{X_{6}},\mathsf{X_{7}}\}$ in Figure 6 as an example. We assume that $\sigma_{\epsilon_{\mathsf{L_{2}}}}=t$, and thus we have $\sigma_{\mathsf{X_{12}},\mathsf{X_{13}}|\mathsf{X_{3}},\mathsf{X_{4}},\mathsf{X_{6}},\mathsf{X_{7}}}=tf_{\mathsf{L_{2}}\rightarrow\mathsf{X_{12}}}f_{\mathsf{L_{2}}\rightarrow\mathsf{X_{13}}}$, and $\sigma_{\mathsf{X_{12}},\mathsf{X_{i}}|\mathsf{X_{3}},\mathsf{X_{4}}}=\sigma_{\mathsf{X_{i}}|(\mathsf{X_{3}},\mathsf{X_{4}})}f_{\mathsf{X_{i}}\rightarrow\mathsf{X_{12}}},6\leq i\leq 7$ and $\sigma_{\mathsf{X_{13}},\mathsf{X_{i}}|\mathsf{X_{3}},\mathsf{X_{4}}}=\sigma_{\mathsf{X_{i}}|(\mathsf{X_{3}},\mathsf{X_{4}})}f_{\mathsf{X_{i}}\rightarrow\mathsf{X_{13}}},6\leq i\leq 7$. We further have $\sigma_{\mathsf{X_{12}},\mathsf{X_{3}}|\mathsf{X_{4}}}=f_{\mathsf{X_{3}}\rightarrow\mathsf{L_{2}}}f_{\mathsf{L_{2}}\rightarrow\mathsf{X_{12}}}+\sum_{i=6}^{7}f_{\mathsf{X_{3}}\rightarrow\mathsf{X_{i}}}f_{\mathsf{X_{i}}\rightarrow\mathsf{X_{12}}}$$\sigma_{\mathsf{X_{13}},\mathsf{X_{3}}|\mathsf{X_{4}}}=f_{\mathsf{X_{3}}\rightarrow\mathsf{L_{2}}}f_{\mathsf{L_{2}}\rightarrow\mathsf{X_{13}}}+\sum_{i=6}^{7}f_{\mathsf{X_{3}}\rightarrow\mathsf{X_{i}}}f_{\mathsf{X_{i}}\rightarrow\mathsf{X_{13}}}$,  and $\sigma_{\mathsf{X_{3}},\mathsf{X_{i}}|\mathsf{X_{4}}}=f_{\mathsf{X_{3}}\rightarrow\mathsf{X_{i}}},6\leq i\leq 7$. We now have $9$ edge coefficients and $9$ equations, and thus by these equations, we can solve these edge coefficients upto $t$. Next we solve $t$ by considering $\mathsf{X_{4}}$. Specifically, we have $f_{\mathsf{X}_{4}\rightarrow\mathsf{X_{i}}}=\sigma_{\mathsf{X}_{4},\mathsf{X_{i}}|\mathsf{X_{3}}}$, where $6\leq i\leq 7$. We also have $f_{\mathsf{X}_{4}\rightarrow\mathsf{L_{2}}}=\sigma_{\mathsf{X}_{4},\mathsf{L_{2}}|\mathsf{X_{3}}}(t)$. By these 3 equations we can solve $f_{\mathsf{X}_{4}\rightarrow\mathsf{L_{2}}}$, $f_{\mathsf{X}_{4}\rightarrow\mathsf{X_{6}}}$, and $f_{\mathsf{X}_{4}\rightarrow\mathsf{X_{7}}}$, upto $t$. Till now all edge coefficients among $\mathsf{X_{3}},\mathsf{X_{4}},\mathsf{L_{2}},\mathsf{X_{6}},\mathsf{X_{7}},\mathsf{X_{12}},\mathsf{X_{13}}$ can be taken as a function of $t$. Finally by the unit variance of $\mathsf{L}$, we further have $t+f_{\mathsf{X_{3}}\rightarrow\mathsf{L_{2}}}^{2}(t)+f_{\mathsf{X_{4}}\rightarrow\mathsf{L_{2}}}^{2}(t)=1$, by which $t$ can be solved. Therefore, all edge coefficients among $\mathsf{X_{3}},\mathsf{X_{4}},\mathsf{L_{2}},\mathsf{X_{6}},\mathsf{X_{7}},\mathsf{X_{12}},\mathsf{X_{13}}$ can be identified.
@@
Example. An example would be in Figure 6. As illustrated in the example in (d), edge coefficients from $\{\mathsf{L_{2}},\mathsf{X_{6}},\mathsf{X_{7}}\}$ to $\{\mathsf{X_{12}},\mathsf{X_{13}}\}$ can be identified. By Lemma 2, $\mathsf{L_{2}}$ can be taken as if it is observed. The reason lies in that other than $\mathsf{X_{12}}$ and $\mathsf{X_{13}}$, there is no descendant of $\{\mathsf{L_{2}},\mathsf{X_{6}},\mathsf{X_{7}}\}$; therefore, from all other variables’ perspective, e.g., $\mathsf{X_{11}}$, $\sigma_{\mathsf{X_{11}},\mathsf{L_{2}}}$ can be calculated as $(\sigma_{\mathsf{X_{11}},\mathsf{X_{12}}}-\Sigma_{i=6}^{7}\sigma_{\mathsf{X_{i}},\mathsf{X_{11}}}f_{\mathsf{X_{i}}\rightarrow\mathsf{X_{12}}})/f_{\mathsf{L_{2}}\rightarrow\mathsf{X_{12}}}$. In other words, $\sigma_{\mathsf{X_{11}},\mathsf{L_{2}}}$ can be calculated using observational data, so $\mathsf{L_{2}}$ can be taken as observed from $\mathsf{X_{11}}$’s perspective. Plus, by the example in (c), we know that all edge coefficients from $\{\mathsf{L_{1}},\mathsf{X_{5}}\}$ to $\{\mathsf{X_{8}},\mathsf{X_{9}},\mathsf{X_{10}}\}$ can be identified and thus similarly by Lemma 2, $\mathsf{L_{1}}$ can be taken as if it can be observed. Finally, we consider edge coefficients among $\mathsf{L_{1}},\mathsf{X_{5}},\mathsf{X_{3}},\mathsf{X_{4}},\mathsf{L_{2}},\mathsf{X_{6}},\mathsf{X_{7}}$. In this situation, $\mathsf{L_{1}}$ and $\mathsf{L_{2}}$ can be taken as if they can be observed and the other variables are themself observed, therefore the edge coefficients among $\mathsf{L_{1}},\mathsf{X_{5}},\mathsf{X_{3}},\mathsf{X_{4}},\mathsf{L_{2}},\mathsf{X_{6}},\mathsf{X_{7}}$ can be easily identified.
A.3 Proof of Theorem 2

### v2

@@ A.3 Proof of Theorem 3
Let $\mathbf{X},\mathbf{Y}$ be two set of variables, we have $\Sigma_{\mathbf{Y}|\mathbf{X}=x}=\Sigma_{\mathbf{Y}}-\Sigma_{\mathbf{Y}\mathbf{X}}\Sigma_{\mathbf{X}}^{-1}\Sigma_{\mathbf{X}\mathbf{Y}}$.
Lemma 2.
Consider a graph $\mathcal{G}$ that satisfies (i) and (ii) in Theorem 3. For an atomic cover $\mathbf{V}$ in $\mathcal{G}$ with one latent variable, $\mathbf{V}=\{\mathsf{L}\}\cup\{\mathsf{X_{i}}\}_{i=1}^{k}$ ($k$ could be zero), if it has an observed pure child $\mathsf{C}$ and the coefficients of the edges from $\mathbf{V}$ to $\mathsf{C}$, i.e., $f_{\mathsf{L}\rightarrow\mathsf{C}}$, $f_{\mathsf{X_{1}}\rightarrow\mathsf{C}}$, …, $f_{\mathsf{X_{k}}\rightarrow\mathsf{C}}$, are known, then for any variable $\mathsf{A}$, such that $\mathsf{A}\neq\mathsf{C}$ and $\mathsf{A}$ is not a descendant of $\mathsf{C}$, $\sigma_{\mathsf{L},\mathsf{A}}$ can be calculated as $(\sigma_{\mathsf{A},\mathsf{C}}-\Sigma_{i=1}^{k}\sigma_{\mathsf{X_{i}},\mathsf{A}}f_{\mathsf{X_{i}}\rightarrow\mathsf{C}})/f_{\mathsf{L}\rightarrow\mathsf{C}}$.
@@
By the definition of atomic covers, all variables in $\mathbf{V}=\{\mathsf{L}\}\cup\{\mathsf{X_{i}}\}_{i=1}^{k}$ are not adjacent. By trek rule and the fact that $\mathsf{C}$ is a pure child of $\mathbf{V}$, all treks from $\mathsf{C}$ to $\mathsf{A}$ go through $\mathbf{V}$, and thus by the trek rule we have $\sigma_{\mathsf{L},\mathsf{A}}f_{\mathsf{L}\rightarrow\mathsf{C}}+\Sigma_{i=1}^{k}\sigma_{\mathsf{X_{i}},\mathsf{A}}f_{\mathsf{X_{i}}\rightarrow\mathsf{C}}=\sigma_{\mathsf{A},\mathsf{C}}$. ∎
Remark: This lemma implies that we can find all edge coefficients of the graph in a bottom-up fashion. Roughly speaking, for a latent variable $\mathsf{L}$ that belongs to an atomic cover $\mathbf{V}$, once we identify $f_{\mathsf{L}\rightarrow\mathsf{C}}$, $f_{\mathsf{X_{1}}\rightarrow\mathsf{C}}$, …, $f_{\mathsf{X_{k}}\rightarrow\mathsf{C}}$ where $\mathsf{C}$ is an observed pure child of $\mathbf{V}$, we can take $\mathsf{L}$ as if it is observed. More specific explanations can be found in the following proof.
See 3
@@
Consider a graph $\mathcal{G}$ that satisfies (i) and (ii) in Theorem 3. We first show that if all the pure children of an atomic cover $\mathbf{V}$ are observed, then all the edge coefficients from the atomic cover to its children are identifiable (up to group sign). To this end, we categorize the scenarios into four cases and prove them separately.
@@
(c) $\mathbf{V}=\{\mathsf{L}\}\cup\{\mathsf{X_{i}}\}_{i=1}^{k}$ ($k\geq 1$) contains a single latent variable and $k$ observed variables, where $\mathbf{V}$ has at least three pure children. We assume that there exist $\mathsf{C_{1}}$, $\mathsf{C_{2}}$, and $\mathsf{C_{3}}$, such that $\mathsf{C_{1}}$, $\mathsf{C_{2}}$, and $\mathsf{C_{3}}$ are pure children of $\mathbf{V}$. Let $\sigma_{\mathsf{L}|\{\mathsf{X_{i}}\}_{i=1}^{k}}=t$. In this case, we have
$\sigma_{\mathsf{C_{1}},\mathsf{C_{2}}|\{\mathsf{X_{i}}\}_{i=1}^{k}}=tf_{\mathsf{L}\rightarrow\mathsf{C_{1}}}f_{\mathsf{L}\rightarrow\mathsf{C_{2}}},$ (8)
$\quad\sigma_{\mathsf{C_{1}},\mathsf{C_{3}}|\{\mathsf{X_{i}}\}_{i=1}^{k}}=tf_{\mathsf{L}\rightarrow\mathsf{C_{1}}}f_{\mathsf{L}\rightarrow\mathsf{C_{3}}},$ (9)
$\sigma_{\mathsf{C_{2}},\mathsf{C_{3}}|\{\mathsf{X_{i}}\}_{i=1}^{k}}=tf_{\mathsf{L}\rightarrow\mathsf{C_{2}}}f_{\mathsf{L}\rightarrow\mathsf{C_{3}}}.$ (10)
By these three equations, we can solve $f_{\mathsf{L}\rightarrow\mathsf{C_{1}}}$,$f_{\mathsf{L}\rightarrow\mathsf{C_{2}}}$, and $f_{\mathsf{L}\rightarrow\mathsf{C_{3}}}$, with the only remaining free parameter $t$. In other words, we have $f_{\mathsf{L}\rightarrow\mathsf{C_{1}}}(t)$,$f_{\mathsf{L}\rightarrow\mathsf{C_{2}}}(t)$, and $f_{\mathsf{L}\rightarrow\mathsf{C_{3}}}(t)$.
Next, we show that $\forall i=1,...,k,j=1,...,3,f_{\mathsf{X_{i}}\rightarrow\mathsf{C_{j}}}$ can be identified. Specifically, as all simple treks between $\mathbf{L}$ and $\mathbf{X}$ contain only observed variables except $\mathbf{L}$, and $\mathbf{L}$ and $\mathbf{X}$ are not directly adjacent, there must exist $\mathbf{\hat{X}}$ such that $\mathbf{\hat{X}}$ d-separates $\mathbf{L}$ and $\mathbf{X}$. Thus, $f_{\mathsf{X_{i}}\rightarrow\mathsf{C_{j}}}\sigma_{\mathsf{X_{i}}|\mathbf{\hat{X}}\cup\mathbf{X}\backslash\mathsf{X_{i}}}=\sigma_{\mathsf{X_{i}}\mathsf{C_{j}}|\mathbf{\hat{X}}\cup\mathbf{X}\backslash\mathsf{X_{i}}}$, by which $f_{\mathsf{X_{i}}\rightarrow\mathsf{C_{j}}}$ can be solved.
@@
Now we solve $t$. The key is that all the edge coefficients along all simple treks between $\mathbf{X}$ and $\mathbf{L}$ can be identified, by using Lemma 2, with only one free parameter $t$. Thus, we can make use of simple trek rule to parameterize $\Sigma_{\mathsf{L},\mathbf{X}}$ as a function of $t$. By Lemma 1, $\sigma_{\mathsf{L}|\mathbf{X}}=t$ can also be formulated as a function of $\Sigma_{\mathsf{L},\mathbf{X}}$, and thus $t$ can be solved. If $\mathbf{V}$ has more than three pure children, we just choose all the combinations of any three pure children.
@@
(d) $\mathbf{V}=\{\mathsf{L}\}\cup\{\mathsf{X_{i}}\}_{i=1}^{k}$ ($k\geq 1$) contains a single latent variable and $k$ observed variables, where $\mathbf{V}$ has two pure children. By Condition 1, there must exist $\mathsf{C_{1}}$, $\mathsf{C_{2}}$ as the pure children of $\mathbf{V}$. If there exists one additional pure child, then it is the same as (c). Plus, if there only exist these two pure children, there must exist $\mathsf{V_{N}}$ such that $\mathsf{V_{N}}$ is a neighbor of $\mathbf{V}$. If $\mathsf{V_{N}}$ is observed, let $\mathsf{X_{N}}=\mathsf{V_{N}}$, otherwise we recursively take $\mathsf{X_{N}}$ as the observed pure children of $\mathsf{V_{N}}$.
Let $\sigma_{\mathsf{L}|\{\mathsf{X_{i}}\}_{i=1}^{k}}=t$. In this case, we have
$\sigma_{\mathsf{C_{1}},\mathsf{C_{2}}|\{\mathsf{X_{i}}\}_{i=1}^{k}}=tf_{\mathsf{L}\rightarrow\mathsf{C_{1}}}f_{\mathsf{L}\rightarrow\mathsf{C_{2}}},$ (11)
$\quad\sigma_{\mathsf{C_{1}},\mathsf{X_{N}}|\{\mathsf{X_{i}}\}_{i=1}^{k}}=tf_{\mathsf{L}\rightarrow\mathsf{C_{1}}}\sigma_{\mathsf{L}\mathsf{X_{N}}},$ (12)
$\quad\sigma_{\mathsf{C_{2}},\mathsf{X_{N}}|\{\mathsf{X_{i}}\}_{i=1}^{k}}=tf_{\mathsf{L}\rightarrow\mathsf{C_{2}}}\sigma_{\mathsf{L}\mathsf{X_{N}}}.$ (13)
Similar to (c), by solving the above, we have $f_{\mathsf{L}\rightarrow\mathsf{C_{1}}}(t)$ and $f_{\mathsf{L}\rightarrow\mathsf{C_{2}}}(t)$.
Next, we show that $\forall i=1,...,k,j=1,...,2,f_{\mathsf{X_{i}}\rightarrow\mathsf{C_{j}}}$ can be identified. Specifically, as all simple treks between $\mathbf{L}$ and $\mathbf{X}$ contain only observed variables except $\mathbf{L}$, and $\mathbf{L}$ and $\mathbf{X}$ are not directly adjacent, there must exist $\mathbf{\hat{X}}$ such that $\mathbf{\hat{X}}$ d-separates $\mathbf{L}$ and $\mathbf{X}$. Thus, $f_{\mathsf{X_{i}}\rightarrow\mathsf{C_{j}}}\sigma_{\mathsf{X_{i}}|\mathbf{\hat{X}}\cup\mathbf{X}\backslash\mathsf{X_{i}}}=\sigma_{\mathsf{X_{i}}\mathsf{C_{j}}|\mathbf{\hat{X}}\cup\mathbf{X}\backslash\mathsf{X_{i}}}$, by which $f_{\mathsf{X_{i}}\rightarrow\mathsf{C_{j}}}$ can be solved.
Now we solve $t$. The key is that all the edge coefficients along all simple treks between $\mathbf{X}$ and $\mathbf{L}$ can be identified, by using Lemma 2, with only one free parameter $t$. Thus, we can make use of simple trek rule to parameterize $\Sigma_{\mathsf{L},\mathbf{X}}$ as a function of $t$. By Lemma 1, $\sigma_{\mathsf{L}|\mathbf{X}}=t$ can also be formulated as a function of $\Sigma_{\mathsf{L},\mathbf{X}}$, and thus $t$ can be solved.
@@
A.4 Proof of Theorem 5
See 5
Proof of Theorem 5.
The proof is a special case of (b) in the proof of Theorem 3. ∎
A.5 Proof of Theorem 4